### v2

# Bringing Generative AI to Adaptive Learning in Education

Abstract
The recent surge in generative AI technologies, such as large language models and diffusion models, have boosted the development of AI applications in various domains, including science, finance, and education. Concurrently, adaptive learning, a concept that has gained substantial interest in the educational sphere, has proven its efficacy in enhancing students’ learning efficiency. In this position paper, we aim to shed light on the intersectional studies of these two methods, which combine generative AI with adaptive learning concepts. By presenting discussions about the benefits, challenges, and potentials in this field, we argue that this union will contribute significantly to the development of the next-stage learning format in education.
Machine Learning, ICML
1 Introduction
Generative AI (GenAI) refers to artificial intelligence models that generate new content, including text, images, audio, and video, based on the patterns and information they have learned from the given training data (Cao et al., 2023). Unlike the traditional Machine Learning (ML) algorithms, which focus on analyzing and interpreting data, GenAI is designed to create new, original outputs and thus able to solve more challenging problems. With the recent success of GenAI technologies, such as large language models and diffusion models, in generating human-like outputs to a wide range of problems in Natural Language Processing (NLP) (Bubeck et al., 2023) and Computer Vision (CV) (Ramesh et al., 2022), the development of applying GenAI to real-world problems is rapidly increasing, which spreads from science (Walters & Murcko, 2020; Lopez et al., 2020), finance (Rane, 2023; Brühl, 2023) to education (Cooper, 2023; Baidoo-Anu & Ansah, 2023).
Meanwhile, Adaptive Learning (AL), which is considered an emerging educational, technological innovation in education (Martin et al., 2020), has been demonstrated with pedagogical benefits, including acceleration, remediation, meta-cognition, mastery-based learning, immediate feedback, and interactive learning (Hattie, 2023). Adaptive learning aims to generate unique learning experiences by accounting for individual differences to improve the scholastic path, learning process, and learner satisfaction in varied learning situations (Liu et al., 2017). As adaptive learning relies on collecting and analyzing data about learners’ interactions, performance, preferences, and progress to tailor the educational experience to their individual needs, ML techniques have been extensively employed in the AL framework for supporting the analysis over large volumes of data about learners’ behavior and assisting in creating personalized learning paths (Gheibi et al., 2021).
Why This Position Paper? The overwhelming performance of GenAI over ML algorithms in challenging data analysis problems provides new opportunities to current ML-based AL systems. By leveraging the advanced capabilities of GenAI in creating human-level responses to problems like reasoning, the new AL systems will be able to finish more complicated AL tasks, such as dynamic learning path planning. Although some pioneering works have started to explore the usage of GenAI for education purposes (Dan et al., 2023; Xiao et al., 2024), there is no systematic review and discussion focusing on bringing GenAI to AL problems. As adaptive learning is increasingly recognized as a significant and promising direction for the future of education (Park & Lee, 2013), in this position paper, we propose to review the related studies and call attention to the opportunities in combining GenAI with AL. Furthermore, by conducting comprehensive discussions on the benefits, challenges, and threats, we hope to shed light on future research in this direction and push adaptive learning into the next chapter.
Our Contributions.
Our position in this paper is that bringing GenAI to AL will present both advantages and challenges to current ML-based AL systems, creating numerous opportunities and original topics for future educational development. To support this claim, we (1) conduct a comprehensive literature review on key ML contributions to adaptive learning, (2) explore the potential benefits and summarize ongoing industrial practices, (3) present in-depth discussions about the challenges and opportunities from the pedagogical perspective, and further extend these considerations to their broader educational impacts.
Related Works.
The impressive performance of GenAI, especially the Large Language Models (LLMs), in generating human-like responses to complicated requests (Bubeck et al., 2023) encourages the recent research in exploring GenAI for education purpose (Dan et al., 2023; Kieser et al., 2023). To provide researchers with a broad overview of the domain, numerous exploratory and survey papers have been proposed. For example, Qadir (2023); Rahman & Watanobe (2023) and Rahman & Watanobe (2023) conclude the applications of ChatGPT to engineering education by analyzing the responses of ChatGPT to the related pedagogical questions. Jeon & Lee (2023) and Mogavi et al. (2023) collect the opinions from different ChatGPT user groups, e.g., educator, learner, researcher, through in-person interviews, online post replies and user logs, and conclude the practical applications of LLMs in education scenarios. Baidoo-Anu & Ansah (2023) and Zhang & Tur (2023) focus on the literature review over the published papers and summarize the progress of the area with structured tables. Although the above works have covered a wide range of existing applications of GenAI in education scenario and provided their long-term visions for future studies, we argue that our position paper is still valuable as none of them focuses on the direction of bringing GenAI to AL. Compared to the general problems such as applying GenAI for education, we propose to present a more concentrated summarization and in-depth discussion toward the adaptive learning topic in this paper.
2 Background
2.1 Adaptive Learning (AL)
Teach-to-the-middle instruction is the most commonly used strategy for transitional teacher-led instructions as it benefits the majority of the students when the teacher resources are limited. However, such method does not fit learners who have different academic ability from the norm. To solve this problem, adaptive learning is developed, which aims to provide an efficient, effective and customised learning experience for students by dynamically adapting learning content to suit their individual abilities or preferences (Park & Lee, 2013; Aleven et al., 2016). In the recent three decades, the effectiveness of AL has been widely recognized by the consistent superior performance reported in different researches’ comparisons between AL users and the comparison sets (Mojarad et al., 2018; Wang et al., 2023b). A valid AL system (Aleven et al., 2016) usually consists of three components: learner module, content module, and instructor module, and its overview is shown in Fig. 1.
Figure 1: An Illustration of Adaptive Learning (AL).
Learner Module, also known as the student model, refers to the learner characteristics of what a student knows and does. The Learner Module includes learner attributes, learner preferences, learner knowledge and proficiency, motivational or emotional aspects of learner behavior, and individual differences that are used to adapt the learning.
Content Module, also known as the expert or domain model, refers to the content or knowledge base for the course. The Content Module could involve concepts that build on each other and include a learning map with relationships between different ideas and how the course content is delivered to the learner.
Instructor Module, also known as the pedagogical model, refers to the algorithm that assists in adapting the instruction based on the content and learner module, and defines what, when, and how adaptation can occur. Some of the adaptation techniques include pacing, the format of instruction, and sequencing. This model provides the base for deciding what content is presented to the learner and can also be called the adaptation model since it describes what is adapted and how it is adapted.
2.2 Generative AI (GenAI)
The origins of Generative AI can be traced back to the mid-1950s with the emergence of the concepts of artificial intelligence and machine learning. Unlike discriminative modeling, which has been the driving force of most progress in artificial intelligence, generative modeling problems are generally more difficult to tackle. Contributing to the rise of deep learning (LeCun et al., 2015), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and Variational Auto-encoders(VAEs) (Kingma & Welling, 2013) first revolutionized image generation and became a cornerstone in the field of GenAI. Moreover, the recent emergence of Transformer (Vaswani et al., 2017), has marked another significant evolution in GenAI within the field of NLP, which has propelled GenAI into a new era. LLMs, such as OpenAI’s GPT series, are the most representative ones of these GenAI models. By taking advantage of the vast number of parameters, extensive pre-training text corpus, and advanced fine-tuning methods (Ouyang et al., 2022), LLMs have shown impressive capabilities in understanding context, generating coherent and contextually relevant responses (Bubeck et al., 2023). Apart from that, further studies (Brown et al., 2020) discover the strong few-shot (in-context) learning ability of LLMs, and such findings encourage the revolution in NLP solutions from task-specific methods to foundation models (Zhou et al., 2023a). Despite the above progress, Generative Diffusion Models (GDMs) (Ho et al., 2020; Song et al., 2020) also bring improvements to the generation quality of images (Dhariwal & Nichol, 2021) and audio signals (Kong et al., 2020). Finally, by connecting GenAI models from different modalities, multi-modal GenAI models are proposed and have achieved promising results in directions like text-to-image (Ramesh et al., 2022), text-to-audio (Popov et al., 2021) and text-to-video generations (Wang et al., 2023a).
3 Machine Learning in Adaptive Learning
Machine learning plays a crucial role in adaptive learning by enabling the system to analyze learners’ behavior, build learning material hierarchies, and provide personalized adaptions to learners. In the following, we summarize three pivotal ML algorithms used by AL systems.
3.1 Knowledge Tracing (KT)
Knowledge proficiency profile is the basis for AL systems and its diagnosing process is well-known as the Knowledge Tracing (KT) problem in literature (Anderson et al., 1990). Generally speaking, the goal of KT is to track the learning progress of students, and use it to instruct the following adjustments toward the students’ learning contents (Abdelrahman et al., 2023). However, as proficiency is not an observable variable, KT methods commonly reformulate the problem as predicting student’s next response in the diagnosis test (Villano, 1992). Specifically, given a student’s past exercising interactions $X$ in a system, the goal of KT is to model student exercise interaction procedure to predict the student’s future performance at timestamp $t+1$, i.e., predicting the probability $p(r_{r+1}=1|e_{t+1},X)$, where $r_{t+1}\in\{0,1\}$ indicates the student’s answer correctness on new exercise $e_{t+1}$. The diagram of knowledge tracing is shown in Fig. 2.
Figure 2: Knowledge Tracing Diagram.
3.1.1 Traditional KT
To achieve that goal, traditional researches developed two series of approaches for KT: Bayesian Knowledge Tracing (BKT) and Factor Analysis Models (FAM). BKT was motivated by the concepts of mastery learning (Corbett, 2000) and commonly uses a probabilistic graphical model, such as Hidden Markov Model (Baker et al., 2008) and Bayesian Belief Network (Villano, 1992) to trace the students’ changing knowledge states as they practice skills. Corbett & Anderson (1994) introduce the first BKT model, which models student’s knowledge state with a binary variable $\{unlearned,learned\}$ and considers transitions from the unlearned state to the learned state. Note that a student may make a mistake while in a learned state or guess correctly in an unlearned state, BKT includes the slip probability $p(S)$ and guess probability $p(G)$ to capture that. In particular, the standard BKT can be expressed as:
$p(r_{n}=1)=p(L_{n})*(1-p(S))+(1-p(L_{n}))*p(G),$
where $r_{n}$ is student’s answer correctness on the $n$-th question and $p(L_{n})$ is the posterior estimation of student’s proficiency, which be estimated by Bayesian formula iteratively:
$p(L_{n})=p(L_{n-1})+(1-p(L_{n-1}))*p(T)$
$p(L_{n-1})=\begin{cases}\frac{p(L_{n-1})*(1-p(S)}{p(L_{n-1})*(1-p(S))+(1-p(L_{n-1}))*p(G)},&r_{n}=1\\ \frac{p(L_{n-1})*p(S)}{p(L_{n-1})*p(S)+(1-p(L_{n-1}))*(1-P(G))},&r_{n}=0\end{cases}$
On the other hand, FAM is theoretically supported by the Item Response Theory (IRT) (Harvey & Hammer, 1999), which estimates student performance by learning a function, usually a logistic function, based on various factors in a population of students who solve a set of problems. The probability $p_{ij}$ that a student $i$ correctly answers a question $j$ is defined as:
$p_{ij}=\mathcal{L}(\theta_{i}-b_{j})=\frac{\exp^{(\theta_{i}-b_{j})}}{1+\exp^{(\theta_{i}-b_{j})}},$
where $b_{j}$ is the difficulty of question $j$ and $\theta_{i}$ is the student $i$’s ability. To estimate $\theta_{i}$, FAM methods take different factors into their consideration. For example, the Additive Factor Model (AFM) (Cen et al., 2008) counts the frequency that a student meets with different skills and uses a linear combination over them to estimate $\theta$. Pavlik Jr et al. (2009) and Thai-Nghe et al. (2012) further add student’s correctness records on the past questions and second order interactions between different skills to increase the complexity of the FAM model for achieving better KT results.
3.1.2 Deep Learning based KT
Inspired by the success of Deep Learning (DL) (LeCun et al., 2015), recent researches on KT have widely adopted deep learning techniques. Deep Knowledge Tracing (DKT) (Piech et al., 2015) was the pioneering study that models KT as a sequence prediction problem and uses Recurrent Neural Network (RNN) (Lipton et al., 2015) to predicate the probability of correctly answering a question at each time step $t$. Furthermore, by taking advantage of the strong fitting capabilities of deep learning methods, recent deep learning based KT models propose a variety of network designs to take more factors into account. For example, memory-augmented KT traces complex Knowledge Concepts (KCs) learned by students and extended DKT by augmenting an external memory structure (Zhang et al., 2017; Abdelrahman & Wang, 2019), attentive KT attempts to incorporate an attention mechanism to mitigates prior model’s limitation in treating all questions in a sequence of interactions equally important (Ghosh et al., 2020; Pandey & Karypis, 2019), graph-based KT leverages graph representation learning techniques (Nakagawa et al., 2019; Tong et al., 2020) to capture such relational structures between questions and their KCs, text-Aware KT leverages the textual features of questions to enhance the performance (Su et al., 2018; Liu et al., 2019). In general, deep learning based KT models have demonstrated their great potential in solving the KT problem.
3.2 Knowledge Concepts Structure Construction
The hierarchical structure of knowledge concepts and learning materials helps AL systems in identifying the appropriate learning contents for students in an efficient manner (Adjei et al., 2016). Because of the heavy manual workload in organizing the huge amount of learning content into a complete knowledge concept framework, following ML methods are widely adopted by AL systems and the diagram is shown as Fig. 3.
Figure 3: Knowledge Concepts Structure Construct Diagram.
3.2.1 Knowledge Concepts Extraction
Knowledge concepts used to be extracted by experts from domain texts. However, the rapid growth of the internet has caused conventional KCs extraction no longer meet the demand for handling large volumes of online data or updating domain knowledge in a timely fashion. To solve the problem, considerable researches aim to develop ML techniques to automate the KCs extraction. The existing ML methods for KCs extraction can be classified into two categories: supervised and unsupervised. The supervised extraction leverages classification models to categorize the text contents into numerous concepts (Firoozeh et al., 2020). However, this technique calls for annotated samples and its performance is heavily relied on the annotation qualities. On the other hand, unsupervised extraction recently dominated the recent KCs extraction because of the appearance of pre-trained language models. The initial unsupervised extraction was built with grammatical rules and semantic rules (Szwed, 2015; Stanković et al., 2016). For example, the Part-of-Speech (POS) tag rules such as Noun+Noun and [Adj or Noun]+Noun are commonly used to discern the KCs. Dictionary and Statistical based methods are later proposed to solve the low recalling issue of the rule-based method. Dictionary based methods (Gong et al., 2017; Levow et al., 2005) utilize a pre-defined concept dictionary to compare words or phrases in the text and extract relevant ones with similarity metrics. Statistical based methods (Boudin, 2018; Bougouin et al., 2013) analyze the word frequency and co-occurrence patterns, and use the rank of statistical features to identify KCs. The semantic-based extraction methods (Xiong et al., 2021; Huang et al., 2023) are the recent proposed ones which utilize the pre-trained models, e.g., word2vec (Mikolov et al., 2013) and BERT (Devlin et al., 2018), to obtain semantic representation of each candidate and apply post-processing filter models to extract target KCs. Because of the deeper meaning and contextual information captured by the pre-trained models, semantic-based extraction methods present the best result in both precision and recall among all unsupervised methods.
3.2.2 Pre-requisite Prediction
The pre-requisite prediction focuses on determining the relationship between two knowledge concepts. To be specific, for a concept pair $(A,B)$, there are four possible relations between them: (1) $A$ is a prerequisite of $B$; (2) $B$ is a prerequisite of $A$; (3) the two concepts are related, but they do not have any prerequisite relation; and (4) the two concepts are unrelated (Talukdar & Cohen, 2012). To simplify the problem, prior studies usually converted this task into a binary classification problem, which can be defined as:
$\mathrm{Preq}(B,A)=\begin{cases}1,A\mathrm{\ is\ a\ prerequisite\ of\ B}\\ 0,\mathrm{otherwise}\end{cases}$
The key components of pre-requisite prediction models are concept representation encoder and classification function. As concepts are usually associated with their description text, the text encoders are commonly employed as concept representation encoder. For the traditional methods, researches (Meng et al., 2016; Gasparetti, 2022) use bag-of-words with word frequencies or TF-IDF weights as concept representations. With the emergence of advanced text representation techniques, such as word2vec (Mikolov et al., 2013) and BERT (Devlin et al., 2018), recent approaches prefer to adopt these pre-trained embedding to capture the high-level semantic similarities between the concept pairs (Bai et al., 2021). The classification function is designed to map the representations of paired concepts to a scalar number which represents the likelihood of the pre-requisite relationship between two concepts. For simplicity, the inputs of classification function are usually the concept representations, but works (Bai et al., 2021) also explore to include manual features like hyperlink count, jaccard similarity and text categories. The common options for classification function include inner product, logistic regression and Multi-Layer Perceptron (MLP). To train pre-requisite prediction models, most studies use the classification-based losses such as binary cross entropy, but there are also studies (Liu et al., 2016) treating the problem as a ranking problem, which leverage the triplet loss.
3.3 Learning Path (LP) Generation
The individual differences cause the uniqueness in each learner’s optimal Learning Path (LP). With knowing learners’ characteristics, including preference, targets, and abilities, AL systems use the following ML methods to generate personalized learning paths, which are shown as Fig. 4.
Figure 4: Learning Path Generation Diagram.
3.3.1 Learner-based LP Generation
Learner-based LP focuses on creating a customized learning path by analyzing characteristics of the learning behavior observed during the learning process. Lin et al. (2013) first develop a personalized innovative learning system with decision tree. It considers students’ demographic information such as age, gender, major and creativity scores during LP recommendation process. Basu et al. (2013) and Salehi & Kamalabadi (2013) take more learner features, such as preference, academic performance, credit requirements, time availability into account, and leverage mixed, weighted, and cascade hybrid methods to form the final recommended learning paths. Salehi et al. (2014) propose a learning material recommendation system framework based on sequential pattern mining and multidimensional attribute-based Collaborative Filtering (CF). Yang & Dong (2017) propose a LP model that allows learning activities and the assessment criteria of their learning outcomes to be explicitly formulated by Bloom’s Taxonomy. Bendahmane et al. (2017) present a competence based approach derived from learning data, learner’s characteristics and their expectations. In this approach, learners were clustered and traced, and finally proper learning paths were presented. Dwivedi et al. (2018) introduce LPRS, which considers learners’ learning styles and knowledge levels and generates a learning path through a variable-length genetic algorithm.
3.3.2 Structure-based LP Generation
As knowledge concepts have strong logical sequences, many LP generation algorithms also exploit the relationship between knowledge elements to provide learners with an efficient learning path to their targets. Chu et al. (2011) propose a LP generation algorithm based on ontology. In this method, the ontology is established based on relationships between knowledge elements, and then a learning path is recommended according to the relationship between them in the ontology. Ouf et al. (2017) propose the framework for smart e-learning ecosystem using ontology and SWRL. Tam et al. (2014) present an explicit semantic analysis, followed by enhancing the ontology analysis through concept clustering, and applied an optimizer to find an optimal learning path of involved concepts or modules. Shmelev et al. (2015) combine the learner’s model with domain knowledge ontology of learning resources for matching, and generate personalized recommendation through the semantic recommendation algorithm. Al-Hassan et al. (2015) propose a hybrid semantic enhanced recommendation approach by combining the new Inferential Ontology-based Semantic Similarity (IOBSS) measure and standard item-based CF approach. Tseng et al. (2017) construct the concept map for adaptive learning and provide educational recommendations for individual students.
4 Bringing GenAI to Adaptive Learning
The recent achievements made by GenAI in the areas such as medicine (Zhang & Kamel Boulos, 2023), programming (Sarkar et al., 2022), and designing (Cai et al., 2023) underscore its exceptional potential in applications compared to the prior machine learning methods. In the following, we will outline the advantages GenAI that can offer to adaptive learning, along with examples of industrial applications. Then, we will provide insights from an educational perspective, delving into the disadvantages and challenges faced by the new framework.
4.1 Benefits and Strengths
4.1.1 Diversity and Dynamics
Integrating generative AI into adaptive learning offers significant benefits, notably its ability to produce dynamic and diverse outputs. Current ML-based adaptive learning systems, despite their automation (Wang et al., 2023b), still fall short of replicating the nuanced interactions of a skilled human teacher. GenAI can bridge this gap by generating real-time, tailored responses and learning plans based on each student’s immediate questions and learning status. Moreover, the recent advancements in large language models for context understanding (Brown et al., 2020) further enhance the potential for providing personalized feedback, offering a substantial leap in optimizing learning outcomes. Apart from providing new features, GenAI holds the potential to enhance current ML models in AL systems. As detailed in Section 3.1, knowledge tracing methods typically use learner’s question practice records, but these are often pre-set and designed for mainstream students, potentially hindering accurate assessment of individual learning status. GenAI, with its dynamic and diverse capabilities, can generate questions tailored to each learner’s past performance, aiding in more effective knowledge tracing. Furthermore, GenAI’s ability to diversify content, such as in math word problems (MWP) (Zhou et al., 2023b), reduces repetition and improves the accuracy of knowledge proficiency assessments. This adaptability makes GenAI a valuable asset in evolving and refining adaptive learning systems.
4.1.2 Multi-modality Convenience
As the strong capabilities of GenAI in dealing with multi-modal data have been verified in multiple studies (Bubeck et al., 2023), applying GenAI will enable adaptive learning to understand student’s learning status from a broader perspective. For example, the current accepted data format to adaptive learning methods are mainly limited to structured text, by integrating GenAI, the AL systems will be able to interpret leaner emotions in speech and facial expressions together with the text input. The enhanced input information will help the following analysis of the student’s learning characteristics such as learning style and learning status (Giannakos & Cukurova, 2023), which will also benefit the precise adaptions to the student’s learning paths. Meanwhile, as GenAI’s outputs can also be multi-modal (Wu et al., 2023), the generated learning materials for knowledge concepts can be more inspiring. For example, with the help of image generation models, the interpretations of geometry problems could be presented with figures or even a short animation, which will be easier for learners to understand compared to prior pure-text responses.
4.1.3 Broad General Knowledges
One of the common challenges for machine learning methods in education is the scarcity of high-quality annotated data, which is caused by its strong privacy properties and high annotation costs. Due to this fact, many of prior models suffer from over-fitting and poor generalization problems. Fortunately, by introducing GenAI, the above issues can be significantly alleviated as GenAI models, especially LLMs, are general-purpose foundation models (Brühl, 2023). Contributing to the broad general knowledge learned from the enormous sized pre-training data, GenAI has been widely recognized as a suitable solution in dealing with data limitation problems (Brown et al., 2020). Additionally, with the recently proposed few-shot learning methods (Dang et al., 2022), the adaptive learning with GenAI is expected to achieve great improvements in performance compared with the prior ones, which are trained from scratches. At last, examples of GenAI models, such as AlphaZero (Silver et al., 2017), and AlphaFold (Senior et al., 2020), in generating unconventional solutions suggest that GenAI may understand and utilize its learned knowledge in ways that are different from our human beings. By bringing GenAI to adaptive learning, we expect to see novel results, e.g., unexplored learning paths, during the usages. With such ”jump out of box” thoughts, we may get chances to explore more improvements to adaptive learning in the future.
4.2 Industrial Practice
Recognizing the benefits, leading institutions in adaptive learning research have begun to experiment with integrating large language models into their existing educational frameworks. In Table 1 of the Appendix, we list some of these pioneering institutions as representative examples of this trend. The majority of these business cases focus on leveraging GenAI to empower their original products and creating personalized learning experience to their customers.
4.3 Disadvantages, Challenges and Potentials
4.3.1 Hallucination
The biggest issue with generative AI currently is its potential to produce content that doesn’t actually exist, a phenomenon known as ”Hallucination” (Ji et al., 2023). Hallucinations happened even in the most advanced generative AI models. Research suggests that educating students about the limitations of using Generative AI is crucial, emphasizing its susceptibility to errors when working with it (Strzelecki & ElArabawy, 2024; Bearman & Ajjawi, 2023). This brings forth numerous opportunities and research questions for the future, such as how to design the adaptive learning system to help students cultivate an appropriate reliance on AI-generated content (Vereschak et al., 2021). Furthermore, hallucination could become a new adaptive teaching method, such as creating new and possibly erroneous educational content and asking students to practice the role of a tutor in correcting AI’s mistakes (Ma et al., 2023).
4.3.2 Capability Decay
The ease and efficiency of obtaining information from GenAI can create a dependency, where students might prefer quick GenAI answers over the process of learning and discovery. This overreliance could erode critical thinking skills and reduce the students’ capacity for self-driven inquiry, as they might become accustomed to receiving immediate answers without engaging in the cognitive effort required for learning. One potential solution is to integrate generative AI as a tool that prompts further questioning and exploration rather than providing definitive answers. Encouraging projects and assignments that require independent research and critical analysis can also help maintain a balance between utilizing generative AI and fostering self-reliance. However, these suggestions only begin to address the broader challenge of integrating generative AI into education without compromising the development of independent learning skills. This issue remains an open question, presenting a wide array of opportunities for educators, technologists, and policymakers to explore innovative solutions that balance the benefits of generative AI with the imperative of nurturing autonomous, curious learners.
4.3.3 Fairness
Fairness is a crucial ethical concern, particularly in education where generative AI technologies like GPT models are being integrated (Fenu et al., 2022). These advancements, while beneficial, bring to light issues of equitable access and unbiased treatment among learners. A notable concern is the varied exposure of learners to different levels of generative AI. For instance, discrepancies between those accessing advanced systems like GPT-4 and those with older versions like GPT-3.5 can lead to unequal knowledge distribution, potentially exacerbating educational gaps. The effective deployment of generative AI in education also hinges on user familiarity and accessibility. Students well-acquainted with AI technologies stand to gain more, while those less experienced may encounter challenges and a steeper learning curve. Introducing standardized training programs might help bridge this gap, though such measures require meticulous implementation and oversight. Additionally, inherent biases within GenAI pose a significant challenge. These systems, limited by the biases in their training data, might offer skewed guidance, making vigilant monitoring essential to ensure fairness. In adaptive learning, where generative AI tailors instructional guidance, the risk of bias amplifies. Personalizing learning experiences is beneficial but needs to be weighed against the risk of unfairness. Ensuring an optimal, equitable, and unbiased educational experience for every learner involves navigating these complexities, making the balance between personalized learning and fairness both a technical hurdle and an ethical necessity in the integration of generative AI into educational systems.
Figure 5: Human-centric Coevolution of GenAI and Education
4.3.4 Coevolution of human, GenAI, and education
The realm of generative AI evolves as an indispensable tool in the educational sphere. This progression, however, underscores a nuanced balance: generative AI’s development must be aligned with, but not overshadow, the broader trajectory of human advancement. Generative AI should serve as an augmentative force, enhancing human capabilities rather than supplanting them. The essence of this coevolution lies in the acknowledgment that generative AI, while a profound enabler, should not assume a dominant or leadership role. Generative AI’s integration into educational systems should amplify human potential without dictating the course of our development. As shown in Fig. 5, we consider the above issues from the following perspectives.
Governance and Ethical Frameworks: How do we establish robust governance structures and ethical frameworks that ensure generative AI’s role in education adheres to human-defined objectives and values? This involves creating policies that prioritize human decision-making in critical educational processes, thereby preventing generative AI from autonomously setting or altering educational agendas.
Human predominance: In what ways can we design generative AI systems that inherently support and enhance human leadership in educational settings? How do we strike a balance between leveraging generative AI’s capabilities for autonomous functioning and maintaining human control over its applications? This balance is crucial to ensure that while generative AI can operate efficiently in certain aspects of education, it does so under the guidelines and constraints set forth by human educators and policymakers.
Educational Goals and generative AI Limitations: In defining educational goals, how to ensure the core of education remains intrinsically human – an arena where human intellect, values, and creativity guide our journey towards a more enlightened future? It’s essential to outline the boundaries of generative AI in handling complex, nuanced aspects of education that are inherently human.
Long-term Impact and Accountability: What mechanisms should we implement to regularly assess the long-term impact of generative AI in education, ensuring that it aligns with human-centric goals and adapting strategies as needed? This involves establishing accountability systems where a golden rule lies on the fact that humans always be the only accountable party that is responsible for the critical decisions, the road map outlined, as well as all outcomes.
5 Discussion
It’s understandable that some readers might have different opinions toward our claims in this paper for a range of reasons. Below, we explore a few of these perspectives and share our discussions in an impartial manner:
Why is our emphasis on adaptive learning rather than other educational perspectives?
Our response to this inquiry can be succinctly encapsulated from two angles: (1) adaptive learning in the educational sphere has a well-established history spanning several decades. Its efficacy in aiding learners to attain enhanced educational outcomes has been consistently validated through numerous studies. By integrating generative AI with adaptive learning, our goal is to maintain and accelerate the evolution of adaptive learning, propelling it into a new era of advancement; (2) the inherently data-driven nature of adaptive learning fosters the widespread adoption of machine learning algorithms within this field. Drawing on insights and methodologies from other ML-centric areas, such as NLP and CV, we are confident that generative AI will serve as a potent catalyst, steering adaptive learning in novel and exciting directions.
What is the scope of this work?
In this paper, we primarily focus on our reviews and discussions of ML-related components in AL systems, particularly because of their close relationship with GenAI. However, just as the example of ”unconventional” solutions advised by GenAI, the deeper explorations in applying GenAI could bring us new visions to educational paradigms. Although this paper might not encompass all these emerging opportunities and challenges, its value remains significant. It aims to illuminate a promising direction for future research and development in this field.
Are all components of adaptive learning compatible with generative AI?
While we’ve highlighted the substantial potential of integrating GenAI into adaptive learning, it’s important to acknowledge the current limitations of GenAI in addressing complex problems. The instructional module of adaptive learning, involving intricate factors like learning theory, pedagogical expertise, and psychological analysis, presents challenges that even domain experts find demanding. Therefore, the next generation of GenAI-driven AL systems will likely still necessitate the nuanced insights of human educators.
How does adaptive learning combined with generative AI broadly impact the field of education?
The integration of GenAI with adaptive learning enhances a previously practical tool, making it more efficient. This advancement not only improves educational methods but also illuminates potential pathways for further revolutions in education technology, particularly concerning universal challenges and risks associated with GenAI models. Moreover, the robust applicability of this method raises the possibility of addressing disparities in educational resources, potentially offering more substantial assistance in this area.
What possible essence can GenAI gain from the education realm?
With the advancement of GenAI models demonstrating human-like proficiency in specific tasks, we can harness learning theories originally devised for humans to enhance GenAI’s efficiency. Recent research supports the viability of this approach (Borges et al., 2023). By applying GenAI in developing sophisticated adaptive learning methods, we gain deeper insights into human learning processes. This knowledge can provide innovative ideas for refining GenAI systems, creating a symbiotic relationship between artificial intelligence development and educational theory.
6 Conclusion
In this paper, we aim to draw attention on the novel research area of integrating generative AI with adaptive learning in education. We believe that the data-centric nature of adaptive learning, which already embraces machine learning, provides an ideal platform for generative AI to enhance the efficacy of existing AL algorithms through its dynamic and diverse output capabilities. However, we also acknowledge the challenges and uncertainties that GenAI introduces to adaptive learning, particularly in sensitive areas like fairness and reliability. Drawing inspiration from GenAI’s advancements in other fields, we remain confident that these concerns will be addressed, transforming these challenges into research opportunities and objectives. While our perspectives might not resonate with every reader, our aim is to spark conversation and encourage exploration in this field. If this paper can inspire future research, then it has successfully served its purpose.
Impact Statements
Our work primarily focuses on advocating for the advancement of educational methods for academic purposes. However, it also brings to light the potential ethical implications of such advancements. Within our discussions, we emphasize crucial issues like fairness, along with the challenges and limitations that accompany these technologies. Although we do not anticipate immediate societal consequences that demand specific attention, we recognize the need for those continuous considerations.
Appendix A Appendix.
In table below, we summarize those pioneering institutions which has begun their industrial practice with bringing generative AI (GenAI) to adaptive learning. The majority of these business cases focus on leveraging GenAI to empower their original products and creating enhanced personalized learning experience to their customers.
Table 1: Industrial practice case studies. USP is the unique selling proposition of the product.
Company USPs Product features
Duolingo Personalized roleplay Uses LLMs to deliver adaptive, highly personalized interactions, closely mimicking real-life conversations. Offers detailed AI-powered post-conversation feedback focusing on the accuracy and appropriateness of responses, identifying strengths and areas for improvement.
Khanmigo 1 on 1 Tutoring Offers a virtual coach, adept at guiding students through problem-solving processes with carefully crafted probing prompts. This method ensures that the learning process is interactive and responsive.
Adaptemy Learning content recommendation Utilizes a triad of models—curriculum, content, and learner to deliver personalized learning experiences, enabling real-time adaptation of content, assessments, and learning paths based on performance.
Squirrel Ai Personalized learning planning Established a large adaptive models (LAM) framework to link the learning contents knowledge graph, student profiles and a study-path planning recommendation system to provide personalized learning paths planning.
DreamBox Learning Math adaptive learning Offers personalized K-8 math learning with supplemental curriculum and a range of live and self-paced resources to meet the learning needs of entire staff.
Chegg 24/7 study support Provides reliable solutions backed by experts, step-by-step explanations and learn tailorings.
Carnegie Learning K–12 education Offers clear solutions family and has been proven to deliver up to 2x performance improvement on standardized tests.
Kahoot! Learning games Provides learning games to make learning fun, easy, and rewarding for everyone.
TAL Education Group Smart learning solutions Offers comprehensive learning services to students from all ages through diversified class formats.